OpenStudy (chrisplusian):

Use polar coordinates to find the volume of the given solid.......Please explain how

OpenStudy (chrisplusian):

Bounded by the paraboloids \[z=7x^2+7y^2 and z=8-x^2-y^2\]

OpenStudy (chrisplusian):

@dumbcow

OpenStudy (dumbcow):

here is how i would do this problem (sorry not polar coord) they intersect at circle x^2 + y^2 = 1 when z = 7 separate volume into lower paraboloid and upper parabaloid each can be found by summing up the areas of the cross-sectional circles lower: \[x^{2} +y^{2} = \frac{z}{7}\] upper: \[x^{2} + y^{2} = 8-z\] \[V = \pi \int\limits_0^7\frac{z}{7} dz + \pi \int\limits_7^8 (8-z) dz\]

OpenStudy (dumbcow):

@chrisplusian